# MATE: LLM-Powered Multi-Agent Translation Environment for Accessibility Applications

### v1

@@ IV-A Evaluation of Interpreter Agents
The performance of several existing LLMs (GLM-4-Flash [14], Llama-3.1-70B-Instruct [27], and GPT-3.5-Turbo [57]) was tested. GLM‑4‑Flash provides free API access, making it widely accessible for zero or low‑budget experimentation. Llama3.1‑70B, as an open‑source model, is relatively cheap for API calls while delivering strong performance, providing an adequate balance between resource consumption and output quality. The GPT models are widely regarded as a high standard in LLM quality, with their robust instruction-following and reliability for general NLP tasks. Hence, GPT-3.5-Turbo was selected as the third model. Importantly, the most advanced closed-source models, such as GPT-4 or Llama3.1‑405B, were avoided because their API usage is expensive, and the marginal performance gains might not justify the extra costs. Thus, our focus was on models that offer strong overall performance while remaining financially sustainable. Three separate interpreter agents were initialized under the configurations of these models, and their performance was evaluated based on the first version of the ModConTT dataset.
[…]
IV-B Evaluation of Task Classification Models
@@
Table III summarizes the performance of a variety of task classification models on the second version of the ModConTT dataset. Within the TF-IDF embeddings, random forest outperformed other ML-based classifiers, achieving an accuracy of 0.650 and an F1-score of 0.629. When using BERT representations, logistic regression achieved the best quality among non-LLM methods with an accuracy of 0.783 and an F1-score of 0.779. However, two of the top-3 performing systems are MAS-based interpreter agents. The GPT-3.5-Turbo agent achieved an accuracy of 0.750 and an F1-score of 0.720, while the fine-tuned BERT model (ModCon-Task-Identifier) agent attained the best overall classification results (accuracy 0.917 and F1-score 0.916), significantly outperforming other tested methods on all the metrics. This performance gap highlights the capacity of LLM-based agents to effectively capture complex task semantics. Hence, the results demonstrate that the use of MAS is justified for the modality conversion problem, as MAS architectures proved to be effective against ML models.
@@
V Limitations and Future Work

### v2

@@ 3.2. ModConTT Dataset Creation
Given the absence of existing datasets appropriate for training and evaluating models in modality conversion task type recognition, we constructed a diverse dataset (ModConTT) by leveraging LLMs to generate prompts aligned with various task types. These tasks include TTS, STT, ITT, ITA, VTT, TTI, ATI, TTV, ATV, and unknown (UNK). Prompts labeled with UNK refer to confusing, unrelated, or irrelevant requests from which the specific prompt type cannot be extracted. The dataset was made in two versions. The first version was created to identify the most accurate of the selected LLMs. These LLMs are available through API calls and are used for interpreter agent initialization. This version consists of 230 samples (20 for each category, except for the UNK class, which has 50 statements). The second version was used to train model-based task type classifiers from scratch. In total, there are 150 observations for the UNK category and 50 unique examples for each of the remaining task types, resulting in 600 prompts overall. Prompts in these two versions do not overlap. In ModConTT, each example includes a user prompt and the correct task type associated with this prompt. Existing large-scale LLMs were used to generate the prompts consistently and diversely so that the data is comprehensive for both training and evaluation. The dataset was designed to be applicable to various modality conversion use cases. Hence, the set of labels is larger than the number of modality adaptation tasks that the MATE system is capable of solving. For instance, MATE does not solve TTV tasks yet; however, TTV examples are present in ModConTT.
3.3. Model Training and Evaluation
In addition to using existing LLMs through API calls, five different classifiers (logistic regression, random forest, SVM, KNN, and CatBoost) were trained on the second version of the ModConTT dataset to identify the most suitable task type recognition model. The objective of these classifiers is to determine the most probable task type $t$ from a set of possible task types $T$ given $P_{user}$. The classification task is defined as:
(4) $\hat{t}=\arg\max_{t\in T}P(\,t\,|\,P_{user};\theta_{\mathcal{M}}),$
where $\theta_{\mathcal{M}}$ represents the learned parameters of a given model $\mathcal{M}$, and $\hat{t}$ is the model prediction. Prompt embeddings were created using the TF-IDF method and the BERT (Devlin et al., 2019) model. Each of the classifiers was trained on both types of embeddings, resulting in 10 models overall. The data set was divided into train and test (90% and 10%, respectively) for all classifiers except CatBoost, where an additional 10% of the train set observations were used as validation. All classifiers were trained with default hyperparameters, ensuring a fair comparison. Besides training classical machine learning models, we fine-tuned a BERT model on our train set. The number of early stopping rounds was set to be 5, and the total number of fine-tuning epochs was 6. The fine-tuned model (ModCon-Task-Identifier) was publicly released on Hugging Face for reproducibility and further research, and all the trained classifiers are available at the project repository. Models were evaluated using standard classification metrics (accuracy, precision, recall, and F1-score). Precision, recall, and F1-score were averaged (using a weighted average strategy) across different classes to get a unique score for each model. In addition, the failure rate (FR) was calculated for LLM-powered interpreter agents, since they are not guaranteed to output a valid label. The failure rate is defined as the number of prompts where an agent failed to output a valid label from the pre-defined set, divided by the total number of prompts processed. Formally, given $N$ total prompts and a set of valid task labels $T$, the failure rate is defined as:
(5) $\text{FR}=\frac{\sum_{i=1}^{N}\mathbb{I}(\mathcal{M}(P_{i})\notin T)}{N},$
where $\mathcal{M}(P_{i})$ is the model’s prediction for prompt $P_{i}$, and $\mathbb{I}(\cdot)$ is the indicator function.
4. Results and Analysis
4.1. Evaluation of Interpreter Agents
@@
The performance of several existing LLMs (GLM-4-Flash (GLM et al., 2024), Llama-3.1-70B-Instruct (Meta AI, 2024), and GPT-3.5-Turbo (Ye et al., 2023)) was tested. GLM‑4‑Flash provides free API access, making it widely accessible for zero or low‑budget experimentation. Llama3.1‑70B, as an open‑source model, is relatively cheap for API calls while delivering strong performance, providing an adequate balance between resource consumption and output quality. The GPT models are widely regarded as a high standard in LLM quality, with their robust instruction-following and reliability for general NLP tasks. Hence, GPT-3.5-Turbo was selected as the third model. Importantly, the most advanced closed-source models, such as GPT-4 or Llama3.1‑405B, were avoided because their API usage is expensive, and the marginal performance gains might not justify the extra costs. Thus, our focus was on models that offer strong overall performance while remaining financially sustainable. Three separate interpreter agents were initialized under the configurations of these models, and their performance was evaluated based on the first version of the ModConTT dataset.
Table 2 shows the comparison of the performance of these three models on the first version of the ModConTT dataset. GPT-3.5-Turbo outperformed both GLM-4-Flash and Llama3.1‑70B on all the metrics, achieving an accuracy of 0.865, precision of 0.880, recall of 0.865, F1-score of 0.865, and a failure rate of 1/230 ($\approx 0.4$%). Hence, the GPT-3.5-Turbo model is recommended for initializing the external LLM-powered interpreter agents. Fig. 2 (d) displays the distribution of prediction failures by task type. The highest failure rate was observed for the UNK class (32%), since this category is the most common in the dataset. The second highest failure rates were detected for STT, ATV, and VTT categories (16% for each one). Tasks involved in converting text and audio files to images were the simplest to recognize, accounting for 20% of the overall failures.
4.2. Evaluation of Task Classification Models
@@
Table 3 summarizes the performance of a variety of task classification models on the second version of the ModConTT dataset. Within the TF-IDF embeddings, random forest outperformed other ML-based classifiers, achieving an accuracy of 0.650 and an F1-score of 0.629. When using BERT representations, logistic regression achieved the best quality among non-LLM methods with an accuracy of 0.783 and an F1-score of 0.779. However, two of the top-3 performing systems are MAS-based interpreter agents. The GPT-3.5-Turbo agent achieved an accuracy of 0.750 and an F1-score of 0.720, while the fine-tuned BERT model (ModCon-Task-Identifier) agent attained the best overall classification results (accuracy 0.917 and F1-score 0.916), significantly outperforming other tested methods on all the metrics. This performance gap highlights the capacity of LLM-based agents to effectively capture complex task semantics. Hence, the results demonstrate that the use of MAS is justified for the modality conversion problem, as MAS architectures proved to be effective against ML models.
@@
5. Limitations and Future Work
5.1. Limitations of the Study